Suppose 0.112 g of zinc bromide is dissolved in 100. mL. of a 27.0 mM aqueous solution of potassium carbonate. Calculate the final molarity of bromide anion in the solution. You can assume the volume of the solution doesn't change when the zinc bromide is dissolved in it. 


Be sure your answer has the correct number of significant digits. 
Answer #1

ZnBr2 + K2CO3 → ZnCO3(s) + 2KBr

moles = (grams / molecular weight)

moles = liters *Molarity

Molecular Weight of ZnBr2 = 65.38 + (2*79.90)

                                               = 225.18 g/mole

moles ZnBr2 = 0.112 g ZnBr2 / 225.18 g/mole

                       = 4.97*10^-4 moles

moles K2CO3 = 0.1 l*0.027 M = 0.0027 moles

Molarity of bromide ion = moles/Volume

                                          = (2*4.97*10^-4)/0.1

                                          = 9.94*10^-3 M